# Bayesian Structure Learning in Undirected Gaussian Graphical Models: Literature Review with Empirical Comparison

### v2

@@ Abstract
Gaussian graphical models provide a powerful framework to reveal the conditional dependency structure between multivariate variables. The process of uncovering the conditional dependency network is known as structure learning. Bayesian methods can measure the uncertainty of conditional relationships and include prior information. However, frequentist methods are often preferred due to the computational burden of the Bayesian approach. Over the last decade, Bayesian methods have seen substantial improvements, with some now capable of generating accurate estimates of graphs up to a thousand variables in mere minutes. Despite these advancements, a comprehensive review or empirical comparison of all recent methods has not been conducted. This paper delves into a wide spectrum of Bayesian approaches used for structure learning and evaluates their efficacy through a simulation study. We also demonstrate how to apply Bayesian structure learning to a real-world data set and provide directions for future research. This study gives an exhaustive overview of this dynamic field for newcomers, practitioners, and experts.
@@ 1 Introduction
One can depict conditional dependencies between a large number of variables using undirected graphical models. Using data to recover the structure of a graphical model is called structure learning. There are many applications of undirected graphical models. In biology, they are used to recover gene networks (Bhadra & Mallick, 2013; Li et al., 2019a; Chandra et al., 2022); in neuroscience, graphical models can discover the connectivity of the brain (Hinne et al., 2014; Dyrba et al., 2020); in economics, they map the relationships between purchases of individual customers (Giudici & Castelo, 2003); in finance, they discover how the credit risks of financial institutions are related (Wang, 2015); and in psychology, they map the relationships between psychological variables (Epskamp et al., 2018; Waldorp & Marsman, 2022).
Within the context of structure learning, Bayesian methods offer several advantages over frequentist approaches. First, Bayesian methods facilitate the integration of prior beliefs, which can potentially enhance the precision of learning the structure. For example, one can have an idea about the risk factors of a disease before seeing any patient data. Second, Bayesian methods estimate the entire probability distribution of parameters, whereas frequentist methods only provide their point estimates. This allows Bayesian methods to incorporate model uncertainty. Despite these advantages, frequentist methods in structure learning are more popular than Bayesian methods due to their computational efficiency and simplicity. However, in the last decade, new Bayesian methods have been proposed that are computationally efficient and accurate, even for large-scale graphs. Meanwhile, the development of several software packages has made Bayesian structure learning accessible and practical; see, for example, the R packages BDgraph (Mohammadi et al., 2022), ssgraph (Mohammadi, 2022), BGGM (Williams & Mulder, 2019) and baygel (Smith et al., 2023a).
@@
In this article, we consider general undirected graphs. Moreover, we assume Gaussian data, i.e., we work with Gaussian graphical models (GGMs). GGMs are the most widely researched undirected graphical models due to their relative simplicity and their variety of applications. The focus of this review is on Bayesian methods in GGMs. However, we do include a short overview of the relevant frequentist methods as they form the starting point for some of their Bayesian counterparts. Moreover, the simulation study contains a frequentist approach to show how Bayesian algorithms perform relative to the better-known frequentist alternative. Our main focus is on the Bayesian methods of the last two decades since these contributed most to the progress in Bayesian structure learning. GGMs form the foundation of several sub-fields that we briefly touch upon too.
We continue in Section 2 with a general introduction to frequentist and Bayesian structure learning in GGMs. Section 3 provides a review of the recent literature, and Section 4 contains a simulation study comparing the performance of seven state-of-the-art Bayesian algorithms and one frequentist algorithm. In Section 5, we apply Bayesian algorithms to a gene data set and discuss typical challenges that arise in real data application of Bayesian structure learning. We end with a short overview of related sub-fields and give future perspectives and recommendations in Section 6.
@@ 2 Structure Learning
This section introduces structure learning. We start with the problem and notation in Subsection 2.1, briefly discuss frequentist approaches in Subsection 2.2, and end with an introduction to Bayesian structure learning in Subsection 2.3.
@@ 2.1 Problem and Notation
Let $(X_{1},X_{2},\dots,X_{p})$ be a vector of random variables, and $X_{-ij}$ denote this vector without the variables $X_{i}$ and $X_{j}$ with $i,j=1,2,\dots,p$ and $i\not=j$. We say that the variables $X_{i}$ and $X_{j}$ are conditionally independent, when
@@
In other words, when the values of the variables $X_{-ij}$ are given, knowing the value of $X_{j}$ does not change the probability distribution of $X_{i}$. When two variables $X_{i}$ and $X_{j}$ are conditionally (in)dependent, they are also referred to as partially (un)correlated.
We aim to estimate the conditional dependencies between any pair of variables in an undirected graph $G=(V,E)$. Here, every node in the node set $V=\{1,\dots,p\}$ corresponds to a random variable, and the edge set $E$ is defined by $E=\{(i,j)\in V\times V:i<j\text{ and }X_{j}\text{ and }X_{i}\text{ are conditionally dependent}\}$. We slightly abuse notation by saying $(i,j)\in G$ when we mean $(i,j)\in E$. We denote the true graph with $G^{*}=(V,E^{*})$ and any other graph with $G$. In general, we do not know the true graph. We only observe a sample of $n$ observations of the variables $X_{1},\dots,X_{p}$. Let $x_{li}$ denote the value of the random variable $X_{i}$ for observation $l$ with $i=1,\dots,p$ and $l=1,\dots,n$. We denote all samples of the $i$th variable as $\mathbf{x}_{i}=(x_{1i},..,x_{ni})^{T}$. Lastly, $\mathbf{X}=(\mathbf{x}_{1},..,\mathbf{x}_{p})$ denotes the $n\times p$ matrix containing all observations of all variables. We assume that all observations are drawn from a multivariate normal distribution with mean $\mathbf{0}$ and an unknown $p\times p$ covariance matrix $\mathbf{\Sigma}^{*}$. That is, $(x_{l1},x_{l2},\dots,x_{lp})\sim\mathcal{N}(\mathbf{0},\mathbf{\Sigma}^{*})$ for all $l=1,\dots,n$.
The problem of finding the graph $G^{*}$ was introduced by Dempster (1972) and coined as covariance selection. It is also referred to as structure learning or graphical model determination. The goal of structure learning is to use the sample $\mathbf{X}$ to infer the properties of $G^{*}$. Let $\mathbf{K}^{*}={\mathbf{\Sigma}^{*}}^{-1}$ with elements $k^{*}_{ij}$, $i,j=1,2,\dots,p$, be the true and unknown precision matrix or concentration matrix. The precision matrix turns out to be a helpful tool in unraveling the structure of the true graph, since
@@
In what follows $\mathbf{K}^{*}$ and $\mathbf{\Sigma}^{*}$ denote the true and unknown precision and covariance matrix, respectively. Likewise, $\mathbf{K}$ and $\mathbf{\Sigma}$ denote any other precision and covariance matrix.
@@ 2.2 Frequentist Structure Learning
This section gives a short overview of frequentist methods and their connections to their Bayesian counterparts. Most frequentist methods in structure learning for GGMs obtain an estimate $\hat{G}=(V,\hat{E})$ of the true graph $G^{*}$ by either solving multiple regressions or employing the penalized likelihood (Liang & Jia, 2023, Chapter 2.2). The regression-based approaches start with the observation that we can express each variable $X_{i}$ in terms of the other variables $X_{-i}$ as follows:
$X_{i}=\sum_{j\not=i}\beta_{ij}X_{j}+\epsilon_{i},\quad i=1,2,\dots,p,$ (3)
where $\epsilon_{i}$ is uncorrelated with $X_{-i}$ and $\beta_{ij}=\frac{k_{ij}^{*}}{k_{ii}^{*}}$ (Lauritzen, 1996). Finding the regression coefficients $\beta_{ij}$ in Equation (3) is thus enough for structure learning, since $\beta_{ij}=0\iff k_{ij}^{*}=0\iff(i,j)\not\in G^{*}$. In their neighbourhood selection method Meinshausen & Bühlmann (2006) use the Lasso regression by Tibshirani (1996) to obtain estimates for all regression coefficients by solving a sequence of optimization problems given by
$\min_{\beta_{ij},i\not=j}||\mathbf{X}_{i}-\sum_{i\not=j}\beta_{ij}\mathbf{X}_{j}||^{2}+\lambda\sum_{i\not=j}|\beta_{ij}|,\quad i=1,\cdots,p,$ (4)
where the first term of the objective function is the residual term and the second term (with the tuning parameter $\lambda\geq 0$) is the $\ell 1$-norm regularization term enabling sparse solutions. Next, they include an edge $(i,j)$, if $\beta_{ij}\not=0$. Similar approaches have been proposed by Yuan (2010), Sun & Zhang (2012), and Liu & Wang (2017), each using different methods to find the coefficients $\beta_{ij}$ in (3). The Bayesian regression approaches discussed in Section 3.3 also use Equation (3), but use Bayesian methods to estimate the regression coefficients.
@@
$\log P(\mathbf{X}|\mathbf{K})=C+\frac{n}{2}\log|\mathbf{K}|+\frac{n}{2}tr(\mathbf{S}\mathbf{K}),$
where $C$ is a constant, $\mathbf{S}=\mathbf{X^{T}}\mathbf{X}$ is the sample covariance matrix, and $tr(\mathbf{A})$ and $|\mathbf{A}|$ are the trace and determinant of a square matrix $\mathbf{A}$, respectively. Penalized likelihood methods aim to find the maximum likelihood estimator (MLE), i.e., the positive definite precision matrix $\mathbf{K}$ that maximizes the log-likelihood. However, this MLE is generally not sparse. Penalized likelihood methods therefore add a penalty function $p(\mathbf{K})$ as a regularization term and minimize the negative penalized log-likelihood:
$\min_{\mathbf{K}\in\mathcal{K}}-\frac{n}{2}\log|\mathbf{K}|-\frac{n}{2}tr(S\mathbf{K})-\lambda p(\mathbf{K}),$ (5)
where $\lambda\geq 0$ is the regularization parameter and $\mathcal{K}$ is the set of all $p\times p$ positive definite matrices. In the case of $\ell_{1}$-norm regularization, the penalty term is given by $p(\mathbf{K})=\sum_{i,j,i\not=j}|k_{ij}|$ and the optimization problem (5) becomes a convex program that Yuan & Lin (2007) solve using the interior point algorithm (Vandenberghe et al., 1998). More efficient, however, is the block coordinate descent approach introduced by Banerjee et al. (2008) and improved by Friedman et al. (2008), who coined it the graphical Lasso (glasso) algorithm. To this day, the glasso algorithm remains the most popular in structure learning and is therefore included in the simulation study in Section 4. Performance bounds of the graphical lasso estimator have been given by Rothman et al. (2008), Lam & Fan (2009), and Ravikumar et al. (2011). Several other penalty functions have been proposed, see, for example, Fan et al. (2009). The tuning parameter $\lambda$ directly impacts the sparsity of the estimate $\mathbf{\hat{K}}$. Several strategies have been proposed for the choice of $\lambda$, including the extended Bayesian information criterion (Foygel & Drton, 2010), the rotation information criterion (Zhao et al., 2012), and $k$-fold cross validation (Bien & Tibshirani, 2011). The Bayesian methods discussed in Section 3.1.3 have a direct connection to these frequentists penalized likelihood approaches: their Bayesian estimate of the precision matrix is equal to the one obtained by solving the minimization problem given by (5).
@@ 2.3 Bayesian Structure Learning
Here, $P(\mathbf{X}|G)$ is the likelihood and denotes the probability of observing the data $\mathbf{X}$ when the true graph is equal to $G$. For each $G$, the prior $P(G)$ describes the given belief that $G$ is the true graph. Lastly, $P(\mathbf{X})$ is the normalizing constant given by $P(\mathbf{X})=\sum_{G\in\mathcal{G}}P(\mathbf{X}|G)P(G)$, where $\mathcal{G}$ denotes the set of all graphs with $p$ nodes. To evaluate the likelihood $P(\mathbf{X}|G)$ in Equation (6), one could use
@@
Often this integral is hard to evaluate. Other Bayesian methods therefore focus on the joint posterior distribution of the true graph and precision matrix given by
@@
$P(\mathbf{X}|\mathbf{K},G)=(2\pi)^{-pn/2}|\mathbf{K}|^{n/2}\exp\left\{\frac{-1}{2}tr(\mathbf{K}\mathbf{S})\right\}.$ (9)
$P(\mathbf{K}|G)$ is the prior distribution on the precision matrix. For each $G$ and $\mathbf{K}$, it describes the belief that when the true graph is equal to $G$, the precision matrix is equal to $\mathbf{K}$.

### v3

@@ Abstract
Gaussian graphical models provide a powerful framework to reveal the conditional dependency structure between multivariate variables. The process of uncovering the conditional dependency network is known as structure learning. Bayesian methods can measure the uncertainty of conditional relationships and include prior information. However, frequentist methods are often preferred due to the computational burden of the Bayesian approach. Over the last decade, Bayesian methods have seen substantial improvements, with some now capable of generating accurate estimates of graphs up to a thousand variables in mere minutes. Despite these advancements, a comprehensive review or empirical comparison of all recent methods has not been conducted. This paper delves into a wide spectrum of Bayesian approaches used for structure learning and evaluates their efficacy through a comprehensive simulation study. We also demonstrate how to apply Bayesian structure learning to a real-world data set and provide directions for future research. This study gives an exhaustive overview of this dynamic field for newcomers, practitioners, and experts.
@@ 1 Introduction
One can depict conditional dependencies between a large number of variables using undirected graphical models. Using data to recover the structure of a graphical model is called structure learning. There are many applications of undirected graphical models. In biology, they are used to recover gene networks (Bhadra & Mallick, 2013; Li et al., 2019a; Chandra et al., 2022); in neuroscience, graphical models illustrate the connectivity of the brain (Hinne et al., 2014; Dyrba et al., 2020); in economics, they map relationships between purchases of customers (Giudici & Castelo, 2003); in finance, they discover how risks of financial institutions are related (Wang, 2015); and, in psychology, they map relationships between psychological variables (Epskamp et al., 2018; Waldorp & Marsman, 2022).
Within structure learning, Bayesian methods offer several advantages over frequentist approaches. First, Bayesian methods integrate prior beliefs, which can enhance the precision of learning the structure. For example, one can have an idea about the risk factors of a disease before seeing any patient data. Second, Bayesian methods estimate the entire probability distribution of parameters, whereas frequentist methods only provide point estimates. This allows Bayesian methods to incorporate model uncertainty. Despite these advantages, frequentist methods in structure learning are more popular than Bayesian methods due to their computational efficiency and simplicity. However, in the last decade, new Bayesian methods were proposed that offer computational efficiency and accuracy, even for large-scale graphs. Meanwhile, the development of several software packages makes Bayesian structure learning accessible and practical; see, for example, the R packages BDgraph (Mohammadi et al., 2022), ssgraph (Mohammadi, 2022), BGGM (Williams & Mulder, 2019) and baygel (Smith et al., 2023a).
@@
In this article, we consider general undirected graphs. Moreover, we assume Gaussian data, i.e., we work with Gaussian graphical models (GGMs). GGMs are the most widely researched undirected graphical models due to their simplicity and variety of applications. The focus of this review is on Bayesian methods in GGMs. Besides, we do include a short overview of relevant frequentist methods as they form the starting point for some of their Bayesian counterparts. Moreover, the simulation study contains a frequentist approach to show how Bayesian methods perform relative to the better-known frequentist alternative. We focus on the Bayesian methods of the last two decades since these contributed most to the progress in Bayesian structure learning. GGMs form the foundation of several sub-fields that we briefly touch upon too.
Section 2 provides a general introduction to frequentist and Bayesian structure learning in GGMs. Section 3 provides a review of the recent literature, and Section 4 contains a simulation study comparing the performance of seven state-of-the-art Bayesian algorithms and one frequentist algorithm. In Section 5, we apply Bayesian methods to a gene data set and discuss typical challenges that arise in real-world applications of Bayesian structure learning. We end with a short overview of related sub-fields and give future perspectives and recommendations in Section 6.
@@ 2 Structure Learning
This section introduces structure learning. We start with the problem and notation in Subsection 2.1, discuss frequentist approaches in Subsection 2.2, and end with an introduction to Bayesian structure learning in Subsection 2.3.
@@ 2.1 Problem and Notation
Let $(X_{1},\dots,X_{p})$ be a vector of random variables, and $X_{-ij}$ denote this vector without the variables $X_{i}$ and $X_{j}$ with $i,j=1,\dots,p$ and $i\not=j$. We say that the variables $X_{i}$ and $X_{j}$ are conditionally independent, when
@@
In other words, when the values of the variables $X_{-ij}$ are given, knowing the value of $X_{j}$ does not change the probability distribution of $X_{i}$. When two variables $X_{i}$ and $X_{j}$ are conditionally (in)dependent, they are also referred to as partially (un)correlated (Lauritzen, 1996).
One can depict the conditional dependencies between any pair of variables in an undirected graph $G=(V,E)$. Here, every node in the node set $V=\{1,\dots,p\}$ corresponds to a random variable, and the edge set $E$ is defined by $E=\{(i,j)\in V\times V:i<j\text{ and }X_{i}\text{ and }X_{j}\text{ are conditionally dependent}\}$. We slightly abuse notation by saying $(i,j)\in G$ when we mean $(i,j)\in E$. We denote the true graph with $G^{*}=(V,E^{*})$. In general, we do not know the true graph. We only observe a sample of $n$ observations of the variables $X_{1},\dots,X_{p}$. Let $x_{li}$ denote the value of the random variable $X_{i}$ for observation $l$ with $i=1,\dots,p$ and $l=1,\dots,n$. We denote all observations of the $i$th variable as $\mathbf{X}_{i}=(x_{1i},\dots,x_{ni})^{T}$. Lastly, $\mathbf{X}=(\mathbf{X}_{1},\dots,\mathbf{X}_{p})$ denotes the $n\times p$ matrix containing all observations of all variables. We assume that all observations are drawn from a multivariate normal distribution with mean $\mathbf{0}$ and an unknown $p\times p$ covariance matrix $\mathbf{\Sigma}^{*}$. That is, $(x_{l1},x_{l2},\dots,x_{lp})\sim\mathcal{N}(\mathbf{0},\mathbf{\Sigma}^{*})$ for all $l=1,\dots,n$.
The problem of finding the graph $G^{*}$ was introduced by Dempster (1972) and coined as covariance selection. It is also referred to as structure learning or graphical model determination. The goal of structure learning is to use the sample $\mathbf{X}$ to infer the properties of $G^{*}$. Let $\mathbf{K}^{*}={\mathbf{\Sigma}^{*}}^{-1}$ with elements $k^{*}_{ij}$, $i,j=1,\dots,p$, be the true and unknown precision matrix or concentration matrix. The precision matrix is a helpful tool in unraveling the structure of the true graph, since
@@ 2.2 Frequentist Structure Learning
This section gives a short overview of frequentist methods and their connections to their Bayesian counterparts. Most frequentist methods in structure learning for GGMs obtain an estimate $\hat{G}=(V,\hat{E})$ of the true graph $G^{*}$ by either solving multiple regressions or employing the penalized likelihood (Liang & Jia, 2024, Chapter 2.2). The regression-based approaches start with the observation that we can express each variable $X_{i}$ in terms of the other variables $X_{-i}$ as follows:
$X_{i}=\sum_{j\not=i}\beta_{ij}X_{j}+\epsilon_{i},\quad i=1,\dots,p,$ (3)
where $\epsilon_{i}$ is uncorrelated with $X_{-i}$ and $\beta_{ij}=\frac{k_{ij}^{*}}{k_{ii}^{*}}$ (Lauritzen, 1996). Finding the regression coefficients $\beta_{ij}$ in Equation 3 is thus enough for structure learning, since $\beta_{ij}=0\iff k_{ij}^{*}=0\iff(i,j)\not\in G^{*}$. In their neighbourhood selection method Meinshausen & Bühlmann (2006) use the Lasso regression by Tibshirani (1996) to obtain estimates for all regression coefficients by solving a sequence of optimization problems given by
$\min_{\beta_{ij},j\not=i}||\mathbf{X}_{i}-\sum_{j\not=i}\beta_{ij}\mathbf{X}_{j}||^{2}+\lambda\sum_{j\not=i}|\beta_{ij}|,\quad i=1,\dots,p,$ (4)
where the first term of the objective function is the residual term and the second term (with the regularization parameter $\lambda\geq 0$) is the $\ell_{1}$-norm regularization term enabling sparse solutions. An edge $(i,j)$ is then included, if $\beta_{ij}\not=0$. Zhao & Yu (2006) show that this approach is consistent, i.e., it recovers the true graph with probability going to $1$ as $n\to\infty$. Similar approaches have been proposed by Yuan (2010), Sun & Zhang (2012), and Liu & Wang (2017), each using different methods to find the coefficients $\beta_{ij}$ in (3). The so-called pseudo-likelihood can be derived from Equation 3 and is used by several Bayesian methods discussed in Sections and 3.1.1 and 3.1.2. Moreover, the Bayesian regression approaches discussed in Section 3.3 also use Equation 3, but use Bayesian methods to estimate the regression coefficients.
@@
$\log P(\mathbf{X}|\mathbf{K})=C+\frac{n}{2}\log|\mathbf{K}|-\frac{n}{2}tr(\mathbf{S}\mathbf{K}),$
where $C$ is a constant, $\mathbf{S}=\frac{1}{n}\mathbf{X^{T}}\mathbf{X}$ is the sample covariance matrix, and $tr(\mathbf{A})$ and $|\mathbf{A}|$ are the trace and determinant of a square matrix $\mathbf{A}$, respectively. Penalized likelihood methods aim to find the maximum likelihood estimator (MLE), i.e., the positive definite precision matrix $\mathbf{K}$ that maximizes the log-likelihood. However, this MLE is generally not sparse. Penalized likelihood methods therefore add a penalty function $p(\mathbf{K})$ as a regularization term and minimize the negative penalized log-likelihood:
$\min_{\mathbf{K}\in\mathcal{K}}-\frac{n}{2}\log|\mathbf{K}|+\frac{n}{2}tr(S\mathbf{K})+\lambda p(\mathbf{K}),$ (5)
where $\mathcal{K}$ is the set of all $p\times p$ positive definite matrices. In the case of $\ell_{1}$-norm regularization, the penalty term is given by $p(\mathbf{K})=\sum_{i,j,i\not=j}|k_{ij}|$ and the optimization problem (5) becomes a convex program that Yuan & Lin (2007) solve using the interior point algorithm (Vandenberghe et al., 1998). More efficient is the block coordinate descent approach introduced by Banerjee et al. (2008) and improved by Friedman et al. (2008), who coined it the graphical Lasso (glasso) algorithm. To this day, the glasso algorithm remains the most popular in structure learning and is therefore included in the simulation study in Section 4. Performance bounds of the graphical lasso estimate $\mathbf{\hat{K}}$ have been proposed in the Frobenius norm (Rothman et al., 2008, Theorem 1), in the element-wise $\ell_{\infty}$ norm (Ravikumar et al., 2011, Corollary 1), spectral norm (Ravikumar et al., 2011, Equation 43b) and element-wise $\ell_{1}$ norm Maathuis et al. (2019, Theorem 14.1.2). Moreover, Ravikumar et al. (2011, Theorem 2) prove that the graph $\hat{G}$ inferred by $\mathbf{\hat{K}}$ converges in probability to the true graph $G^{*}$. Several other penalty functions $p(\mathbf{K})$ have been proposed, see, for example, Fan et al. (2009). The tuning parameter $\lambda$ directly impacts the sparsity of the estimate $\mathbf{\hat{K}}$. Several strategies have been proposed for the choice of $\lambda$, including the extended Bayesian information criterion (Foygel & Drton, 2010), the rotation information criterion (Zhao et al., 2012), and $k$-fold cross validation (Bien & Tibshirani, 2011). The Bayesian methods discussed in Section 3.1.3 have a direct connection to these frequentists penalized likelihood approaches: their Bayesian estimate of the precision matrix is equal to the one obtained by solving the minimization problem given by (5).
@@ 2.3 Bayesian Structure Learning
Here, $P(\mathbf{X}|G)$ is the likelihood and denotes the probability of observing the data $\mathbf{X}$ when the true graph is equal to $G$. For each $G$, the prior $P(G)$ describes the given belief that $G$ is the true graph. Lastly, $P(\mathbf{X})$ is the normalizing constant given by $P(\mathbf{X})=\sum_{G\in\mathcal{G}}P(\mathbf{X}|G)P(G)$, where $\mathcal{G}$ denotes the set of all graphs with $p$ nodes. To evaluate the likelihood $P(\mathbf{X}|G)$ in Equation 6, one could use
@@
$P(\mathbf{K}|G)$ is the prior distribution on the precision matrix. For each $G$ and $\mathbf{K}$, it describes the belief that when the true graph is equal to $G$, the precision matrix is equal to $\mathbf{K}$. Often the integral in (7) is hard to evaluate. Other Bayesian methods therefore focus on the joint posterior distribution of the true graph and precision matrix given by
@@
$P(\mathbf{X}|\mathbf{K},G)=(2\pi)^{-pn/2}|\mathbf{K}|^{n/2}\exp\left\{-\frac{n}{2}tr(\mathbf{K}\mathbf{S})\right\}.$ (9)